# The Equational Theories Project: Advancing Collaborative Mathematical Research at Scale

### v1

@@ 4.3. The project template
When working on a formalization project, there are many moving pieces that need to work in concert. At the core level, there is the project set up by Lean’s build and dependency management system lake. But in addition to that, there are several pieces, including the aforementioned blueprint tool, as well as scripts that a user may choose to run to visualise various aspects of the project, or check the project in specific ways, or compile documentation automatically as the project advances. These additional tasks are accomplished by a number of external tools, and combining them in a mutually compatible way can be challenging. We side-stepped most of these issues by using the GitHub template repository of Pietro Monticone [Monticone_LeanProject_2025]. At the same time, when we began the project, the template in place was suited for more conventional formalization projects and the tooling they required. It also did not include the scripts that enabled automated project management support that we added, as well as support for deploying our visualisation tools and the paper. Over the course of the project, the leanproject template in turn received substantial new additions. One elementary example is the addition of git pre-push hooks, which are scripts that perform a basic sanity check on the local working copy of a contributor before pushing their contributions to the central GitHub repository.
@@ 4.4. The Lean Zulip chat forum
The Lean community traditionally congregates on the leanprover Zulip chat forum888leanprover.zulipchat.com. Our project was coordinated and organised primarily from this forum. At the beginning we created a channel called Equational. Zulip allows the creation and management of discussion topics within the scope of a channel. We made extensive use of the Zulip channel for several purposes. In the beginning it became the gathering point for new contributors. The new contributions process was designed and discussed on this forum. Later, topics were created for each specific technical topic, including the metatheory and its formalization, specific design decisions, specific implications and anti-implications, design of tools, etc. As shown in Figure˜7, the Zulip chat served as the beginning of the contributions process for each piece of the project. Contributors first discussed their proposed contributions or specific problems they tackled on Zulip before following the steps of claiming tasks on GitHub, writing a blueprint write up and/or formalization.
@@ 4.5. Organising the collaboration: the precedent set by the PFR project
When five people collaborate in person, splitting up the research on a question into subtasks and assigning them to collaborators can be accomplished by discussion and consensus. When there are more than fifty collaborators working together online, a more systematic approach is required. In previous formalization projects such as the formalization of the proof of the Polynomial Freiman–Ruzsa (PFR) conjecture [PFR_Tao_Dilles_2023], tasks were managed over the Lean zulipchat forum. The organiser of the project, Terence Tao, posted a series of message threads. Each thread corresponded to a list of outstanding tasks. These tasks were then claimed by collaborators on Zulip. The claims were recorded on a first-come first-served basis by the organiser by tagging the respective users against the tasks. Contributors could claim any open task and disclaim tasks if they couldn’t finish it, with the organiser keeping track of these requests. This system allowed contributors to take their time to flesh out their work, without worrying about competing claims to the same task. Further, it helped the organisers track the task assignment and communicate with the respective collaborators to track and ascertain progress. Unfortunately, this involved a lot of manual and time-consuming management of the task list by organisers. In this project, we automated several pieces of this approach. This freed up organisers to help contributors and review their contributions.
@@ 4.6. Organizing the collaboration in this project
We adopted tools that are familiar to software engineers as ticket systems but are also known in the wider world of industrial production, such as the kanban system. Our project dashboard was built using the GitHub projects feature. We were able to encode some pieces of our automation using the standard GitHub-provided interface. For the rest, we relied on continuous integration scripts (hereon CI). The exact flow of contributions is specified in the CONTRIBUTING.md file of the project repository [The_Equational_Theories_repository]. Briefly,
@@
    Upon completion of the task, the pull request received reviews, while the CI automatically checked that the project compiled and passed additional checks such as Lean’s environment replay tool leanchecker and the semi-external checker lean4lean [lean4lean].
@@
We note that our method has since been adopted by other major formalization projects including the one to formalize Fermat’s Last Theorem [FLT_Lean].
@@
Figure 8. A snapshot of the project dashboard as of July 30, 2025
@@ 4.7. Trusting ITPs to scale collaboration
In our project we used the interactive theorem prover (hereon ITP) Lean 4 [the_lean4_paper] precisely to address these issues of scaling. The contents of this section are common knowledge in the ITP and ITP-adjacent research communities. The exposition is intended to be useful to a user of ITPs.
@@
  • Ideally the environment only consists of definitions and lemmata already checked by the kernel. Thus their correct usage in the proof of subsequent theorems is valid. However this assumption is often not entirely true in practical implementations of theorem provers for efficiency reasons. In this case, trust in the kernel is restored by replaying each assumption from scratch in the kernel. This can be accomplished by external checkers. Such checkers, which are ideally independent implementations of the kernel, can also guard against implementation mistakes in the kernel. In this project we used the environment replay tool lean4checker and the lean4lean [lean4lean] tool. To our knowledge this was the first such use of the lean4lean tool.
@@
Users of Lean only interact directly with the higher level programming language and usually get confirmation of the correctness of their proofs through the editor interface. Further, collaborators on a project such as ours are likely to deploy automated tools such as SAT solvers, SMT solvers, first-order theorem provers, and perhaps even modern AI tools. These tools usually produce proof certificates which are imported or inserted into a Lean source file. Some modern AI tools are integrated into code editors which might automatically produce or edit even the statements of theorems and the definitions they deal with. Given the limits to trust mentioned above, a productive and useful collaboration using Lean also requires a collaboration and verification infrastructure combining human effort and automated tools. It is in this context that we discuss the project infrastructure. It is a concrete answer to the questions posed by one of the authors at the beginning of this project [Tao_blog_Sep_2024] that combines tools from the ITP community and the software engineering community. All these tools already exist. The goal of this exposition is to explain how they address the concerns described above.
@@ 5.1. Finite magmas
A finite magma $\mathcal{M}$} of size $n$ can be assumed without loss of generality to have carrier $\{1,\dots,n\}$ and described by specifying the multiplication table $\diamond\colon\{1,\dots,n\}\times\{1,\dots,n\}\to\{1,\dots,n\}$. By generating a list of all the equational laws $\mathrm{E}j$, $j=1,\dots,4694$ satisfied by this magma, one can create refutations: if ${\mathcal{M}}\models\mathrm{E}j$ and ${\mathcal{M}}\not\models\mathrm{E}k$, then clearly $\mathrm{E}j\not\models_{\mathrm{fin}}\mathrm{E}k$ and hence also $\mathrm{E}j\not\models\mathrm{E}k$. (As mentioned previously, these statements were organized in Lean using the Facts statement.) It is feasible to brute force over all $\sum_{n=2}^{4}n^{n^{2}}\approx 4.3\times 10^{9}$ non-trivial magmas of size at most $4$ to obtain many refutations of this type. By performing brute force over all magmas up to size $4$, a total of 13 632 56613\,632\,566 implications ($61.9\%$ of all implications, and $96.3\%$ of the false ones) can be refuted with $524$ distinct magmas. Of these implications, 13 345 05313\,345\,053 were refuted with magmas of size [^9] For an earlier computer exploration of size $3$ magmas, see [berman-burris]. $3$, with the remaining 415 293415\,293 requiring magmas of size $4$. Performing this search took 165 CPU-hours.
@@
It is a result of Kisielewicz [Kisielewicz] that every law $\mathrm{E}n$ with $n\leq 4694$ is either equivalent to the singleton law $\mathrm{E}2$, or else has a non-trivial finite model; in other words, the implications $\mathrm{E}n\models\mathrm{E}2$ and $\mathrm{E}n\models_{\mathrm{fin}}\mathrm{E}2$ are equivalent for $n\leq 4694$. Our brute force search revealed that in the latter case there is always a model of size $2\leq n\leq 5$, with the lone exception of $\mathrm{E}1286$ (and its dual $\mathrm{E}2301$), for which the smallest non-trivial finite model was of size $7$, as presented in Example˜5.2 below. In fact, most of the $4694$ laws either only had trivial models, or had a size $2$ model, as shown in Table˜1.
@@
Size of smallest non-trivial model Number of laws
Trivial only $1496$
22 $3136$
33 $32$
44 $14$
55 $14$
77 22
@@ 55 $14$
The laws satisfied by a given finite magma $\mathcal{M}$} need not be finitely axiomatizable. The smallest example is the three-element magma $\{0,1,2\}$ with $1\diamond 2=1$, $2\diamond 1=2\diamond 2=2$, and $x\diamond y=0$ for all other $x,y$ [murskii-1]. It was also shown in [murskii-2] that “almost all” magmas $\mathcal{M}$} (in a certain precise sense) are idemprimal [^10] A magma is idemprimal if every idempotent function is expressed by a term. This is weaker than being primal (in which every function is expressible by a term), but stronger than being quasiprimal (in which the discriminator $f(a,b,c)$, defined to equal $c$ if $a=b$ and $a$ otherwise, is expressed by a term); quasiprimality is sufficient to show that all other finite magmas satisfying the laws of $\mathcal{M}$} are isomorphic to products of submagmas of $\mathcal{M}$}. By combining these observations with [Padmanabhan], one can also show that quasiprimal finite magmas can be axiomatized by a single equation. We thank Stanley Burris for these comments, as well as the other references and observations in this remark., which implies that their laws are finitely axiomatizable, and all other finite magmas satisfying these laws are isomorphic to powers of $\mathcal{M}$}.
@@ 5.2. Linear models
for all rings $M$. For commutative rings, this criterion can be checked in an automatable fashion by standard Gröbner basis techniques; in the noncommutative case one can use methods such as the diamond lemma [diamond-lemma].
@@
As essentially observed in [austin], if there is a commutative linear counterexample to an implication $\mathrm{E}\models{\mathrm{E}^{\prime}}$, then by the Lefschetz principle this counterexample can be realized in a finite field ${\mathbb{F}}_{q}$ for some prime power $q$ (and by the Chebotarev density theorem one can in fact take $q$ to be a prime, so that the carrier is of the form $\mathbb{Z}/p\mathbb{Z}$ for some prime $p$), so that one also has $\mathrm{E}\models_{\mathrm{fin}}{\mathrm{E}^{\prime}}$. As such, we have found that an effective way to refute implications by the commutative linear magma method is to simply perform a brute force search over linear magmas $x\diamond y=ax+by$ in $\mathbb{Z}/p\mathbb{Z}$ for various triples $(p,a,b)$.
@@ 5.5. Greedy constructions
We have found greedy extension methods, or greedy methods for short, to be a powerful way to refute implications, especially when the carrier $M$ is allowed to be infinite. Such constructions have a long history in model theory, with possibly the earliest [^11] We thank Stanley Burris for this reference. such construction due to Skolem [skolem]. A basic implementation of this method is as follows. To build a magma operation $\diamond\colon M\times M\to M$ that satisfies one law $\mathrm{E}$ but not another $\mathrm{E}^{\prime}$}, one can first consider partial magma operations $\diamond\colon\Omega\to M$, defined on some subset $\Omega$ of $M\times M$. Thus $x\diamond y$ is defined if and only if $(x,y)\in\Omega$. A magma operation is then simply a partial operation which is total in the sense that $\Omega=M\times M$. We say that a partial magma operation is finitely supported if $\Omega$ is finite.
@@ 5.6. Modifying base models
Since our goal is simply to find a finite model — and any candidate can be checked directly — we are not limited to transformations that can be rigorously justified. SAT solvers like Glucose [DBLP:conf/ijcai/AudemardS09, DBLP:conf/cp/AudemardS12] inspired by the earlier MiniSAT [DBLP:conf/sat/EenS03], via convenient interfaces like PySAT [imms-sat18, itk-sat24], other fast SAT solvers such as Kissat [BiereFazekasFleuryHeisinger-SAT-Competition-2020-solvers], counterexample finders like Mace4 [prover9-mace4], and more general ATPs like Vampire [DBLP:conf/cav/KovacsV13] which can be used as solvers, all succeeded in finding useful magmas following this approach.
@@ 6.3. Canonization
As a conclusion for this section, we note that a very general strategy for building canonizers comes from the theory of rewrite systems, see e.g. Baader & Nipkow [term-rewriting]. In that setting one defines rewriting as a transformation on words (or terms), and if this transformation is terminating and confluent (intuitively, rewrites cannot go on forever, or diverge forever), then one may simply pick the transformation which sends a term to its normal form as a canonizer.
@@ 6.4. Unique factorization
If (i) holds, then we have $\mathrm{x}\diamond\mathrm{y}=\mathrm{x}$ must hold in ${\mathcal{M}}_{X}/\sim_{\mathrm{E}}$, hence $\mathrm{E}854$ would imply $\mathrm{E}4$. However, it is possible to refute this implication by a finite counterexample.
Similarly, if (iii) held, then $\mathrm{E}854$ would have to imply $\mathrm{E}10$, but this can also be refuted by a finite magma.

### v2

@@ 4.3. The project template
When working on a formalization project, there are many moving pieces that need to work in concert. At the core level, there is the project set up by Lean’s build and dependency management system lake. But in addition to that, there are several pieces, including the aforementioned blueprint tool, as well as scripts that a user may choose to run to visualise various aspects of the project, or check the project in specific ways, or compile documentation automatically as the project advances. These additional tasks are accomplished by a number of external tools, and combining them in a mutually compatible way can be challenging. We side-stepped most of these issues by using the GitHub template repository of Pietro Monticone [52]. At the same time, when we began the project, the template in place was suited for more conventional formalization projects and the tooling they required. It also did not include the scripts that enabled automated project management support that we added, as well as support for deploying our visualisation tools and the paper. Over the course of the project, the leanproject template in turn received substantial new additions. One elementary example is the addition of git pre-push hooks, which are scripts that perform a basic sanity check on the local working copy of a contributor before pushing their contributions to the central GitHub repository.
@@ 4.4. The Lean Zulip chat forum
The Lean community traditionally congregates on the leanprover Zulip chat forum [^6] https://leanprover.zulipchat.com. Our project was coordinated and organised primarily from this forum. At the beginning we created a channel called Equational. Zulip allows the creation and management of discussion topics within the scope of a channel. We made extensive use of the Zulip channel for several purposes. In the beginning it became the gathering point for new contributors. The new contributions process was designed and discussed on this forum. Later, topics were created for each specific technical topic, including the metatheory and its formalization, specific design decisions, specific implications and anti-implications, design of tools, etc. As shown in Figure˜7, the Zulip chat served as the beginning of the contributions process for each piece of the project. Contributors first discussed their proposed contributions or specific problems they tackled on Zulip before following the steps of claiming tasks on GitHub, writing a blueprint write up and/or formalization.
@@ 4.5. Organising the collaboration: the precedent set by the PFR project
When five people collaborate in person, splitting up the research on a question into subtasks and assigning them to collaborators can be accomplished by discussion and consensus. When there are more than fifty collaborators working together online, a more systematic approach is required. In previous formalization projects such as the formalization of the proof of the Polynomial Freiman–Ruzsa (PFR) conjecture [22], tasks were managed over the Lean zulipchat forum. The organiser of the project, Terence Tao, posted a series of message threads. Each thread corresponded to a list of outstanding tasks. These tasks were then claimed by collaborators on Zulip. The claims were recorded on a first-come first-served basis by the organiser by tagging the respective users against the tasks. Contributors could claim any open task and disclaim tasks if they couldn’t finish it, with the organiser keeping track of these requests. This system allowed contributors to take their time to flesh out their work, without worrying about competing claims to the same task. Further, it helped the organisers track the task assignment and communicate with the respective collaborators to track and ascertain progress. Unfortunately, this involved a lot of manual and time-consuming management of the task list by organisers. In this project, we automated several pieces of this approach. This freed up organisers to help contributors and review their contributions.
@@ 4.6. Organizing the collaboration in this project
We adopted tools that are familiar to software engineers as ticket systems but are also known in the wider world of industrial production, such as the kanban system. Our project dashboard was built using the GitHub projects feature. We were able to encode some pieces of our automation using the standard GitHub-provided interface. For the rest, we relied on continuous integration scripts (hereon CI). The exact flow of contributions is specified in the CONTRIBUTING.md file of the project repository [13]. Briefly,
@@
    Upon completion of the task, the pull request received reviews, while the CI automatically checked that the project compiled and passed additional checks such as Lean’s environment replay tool leanchecker and the semi-external checker lean4lean [16].
@@
We note that our method has since been adopted by other major formalization projects including the one to formalize Fermat’s Last Theorem [15].
@@
Figure 8. A snapshot of the project dashboard as of 30 July 2025
@@ 4.7. Trusting ITPs to scale collaboration
In our project we used the interactive theorem prover (hereon ITP) Lean 4 [53] precisely to address these issues of scaling. The contents of this section are common knowledge in the ITP and ITP-adjacent research communities. The exposition is intended to be useful to a user of ITPs.
@@
  • Ideally the environment only consists of definitions and lemmata already checked by the kernel. Thus their correct usage in the proof of subsequent theorems is valid. However this assumption is often not entirely true in practical implementations of theorem provers for efficiency reasons. In this case, trust in the kernel is restored by replaying each assumption from scratch in the kernel. This can be accomplished by external checkers. Such checkers, which are ideally independent implementations of the kernel, can also guard against implementation mistakes in the kernel. In this project we used the environment replay tool lean4checker and the lean4lean [16] tool. To our knowledge this was the first such use of the lean4lean tool.
@@
Users of Lean only interact directly with the higher level programming language and usually get confirmation of the correctness of their proofs through the editor interface. Further, collaborators on a project such as ours are likely to deploy automated tools such as SAT solvers, SMT solvers, first-order theorem provers, and perhaps even modern AI tools. These tools usually produce proof certificates which are imported or inserted into a Lean source file. Some modern AI tools are integrated into code editors which might automatically produce or edit even the statements of theorems and the definitions they deal with. Given the limits to trust mentioned above, a productive and useful collaboration using Lean also requires a collaboration and verification infrastructure combining human effort and automated tools. It is in this context that we discuss the project infrastructure. It is a concrete answer to the questions posed by one of the authors at the beginning of this project [65] that combines tools from the ITP community and the software engineering community. All these tools already exist. The goal of this exposition is to explain how they address the concerns described above.
@@ 5.1. Finite magmas
A finite magma $\mathcal{M}$} of size $n$ can be assumed without loss of generality to have carrier $\{1,\dots,n\}$ and described by specifying the multiplication table $\diamond\colon\{1,\dots,n\}\times\{1,\dots,n\}\to\{1,\dots,n\}$. By generating a list of all the equational laws $\mathrm{E}j$, $j=1,\dots,4694$ satisfied by this magma, one can create refutations: if ${\mathcal{M}}\models\mathrm{E}j$ and ${\mathcal{M}}\not\models\mathrm{E}k$, then clearly $\mathrm{E}j\not\models_{\mathrm{fin}}\mathrm{E}k$ and hence also $\mathrm{E}j\not\models\mathrm{E}k$. (As mentioned previously, these statements were organized in Lean using the Facts statement.) It is feasible to brute force over all $\sum_{n=2}^{4}n^{n^{2}}\approx 4.3\times 10^{9}$ non-trivial magmas of size at most $4$ to obtain many refutations of this type. By performing brute force over all magmas up to size $4$, a total of 13 632 56613\,632\,566 implications ($61.9\%$ of all implications, and $96.3\%$ of the false ones) can be refuted with $524$ distinct magmas. Of these implications, 13 345 05313\,345\,053 were refuted with magmas of size [^7] For an earlier computer exploration of size $3$ magmas, see [11]. $3$, with the remaining 415 293415\,293 requiring magmas of size $4$. Performing this search took 165 CPU-hours.
@@
It is a result of Kisielewicz [33] that every law $\mathrm{E}n$ with $n\leq 4694$ is either equivalent to the singleton law $\mathrm{E}2$, or else has a non-trivial finite model; in other words, the implications $\mathrm{E}n\models\mathrm{E}2$ and $\mathrm{E}n\models_{\mathrm{fin}}\mathrm{E}2$ are equivalent for $n\leq 4694$. Our brute force search revealed that in the latter case there is always a model of size $2\leq n\leq 5$, with the lone exception of $\mathrm{E}1286$ (and its dual $\mathrm{E}2301$), for which the smallest non-trivial finite model was of size $7$, as presented in Example˜5.2 below. In fact, most of the $4694$ laws either only had trivial models, or had a size $2$ model, as shown in Table˜1.
@@
Size of smallest non-trivial model 22 33 44 55 77 Trivial only
Number of laws of order 0 11 11
  Number of laws of order 11 33 22
  Number of laws of order 22 $27$ $12$
  Number of laws of order 33 $229$ $14$ 22 $119$
  Number of laws of order 44 $2876$ $18$ $12$ $14$ 22 $1362$
        total (order $\leq 4$) $3136$ $32$ $14$ $14$ 22 $1496$
@@
The laws satisfied by a given finite magma $\mathcal{M}$} need not be finitely axiomatizable. The smallest example is the three-element magma $\{0,1,2\}$ with $1\diamond 2=1$, $2\diamond 1=2\diamond 2=2$, and $x\diamond y=0$ for all other $x,y$ [54]. It was also shown in [55] that “almost all” magmas $\mathcal{M}$} (in a certain precise sense) are idemprimal [^8] A magma is idemprimal if every idempotent function is expressed by a term. This is weaker than being primal (in which every function is expressible by a term), but stronger than being quasiprimal (in which the discriminator $f(a,b,c)$, defined to equal $c$ if $a=b$ and $a$ otherwise, is expressed by a term); quasiprimality is sufficient to show that all other finite magmas satisfying the laws of $\mathcal{M}$} are isomorphic to products of submagmas of $\mathcal{M}$}. By combining these observations with [58], one can also show that quasiprimal finite magmas can be axiomatized by a single equation. We thank Stanley Burris for these comments, as well as the other references and observations in this remark., which implies that their laws are finitely axiomatizable, and all other finite magmas satisfying these laws are isomorphic to powers of $\mathcal{M}$}.
@@ 5.2. Linear models
for all rings $M$. For commutative rings, this criterion can be checked in an automatable fashion by standard Gröbner basis techniques; in the noncommutative case one can use methods such as the diamond lemma [10].
@@
As essentially observed in [5], if there is a commutative linear counterexample to an implication $\mathrm{E}\models{\mathrm{E}^{\prime}}$, then by the Lefschetz principle this counterexample can be realized in a finite field ${\mathbb{F}}_{q}$ for some prime power $q$ (and by the Chebotarev density theorem one can in fact take $q$ to be a prime, so that the carrier is of the form $\mathbb{Z}/p\mathbb{Z}$ for some prime $p$), so that one also has $\mathrm{E}\models_{\mathrm{fin}}{\mathrm{E}^{\prime}}$. As such, we have found that an effective way to refute implications by the commutative linear magma method is to simply perform a brute force search over linear magmas $x\diamond y=ax+by$ in $\mathbb{Z}/p\mathbb{Z}$ for various triples $(p,a,b)$.
@@ 5.5. Greedy constructions
We have found greedy extension methods, or greedy methods for short, to be a powerful way to refute implications, especially when the carrier $M$ is allowed to be infinite. Such constructions have a long history in model theory, with possibly the earliest [^9] We thank Stanley Burris for this reference. such construction due to Skolem [63]. A basic implementation of this method is as follows. To build a magma operation $\diamond\colon M\times M\to M$ that satisfies one law $\mathrm{E}$ but not another $\mathrm{E}^{\prime}$}, one can first consider partial magma operations $\diamond\colon\Omega\to M$, defined on some subset $\Omega$ of $M\times M$. Thus $x\diamond y$ is defined if and only if $(x,y)\in\Omega$. A magma operation is then simply a partial operation which is total in the sense that $\Omega=M\times M$. We say that a partial magma operation is finitely supported if $\Omega$ is finite.
@@ 5.6. Modifying base models
Since our goal is simply to find a finite model — and any candidate can be checked directly — we are not limited to transformations that can be rigorously justified. SAT solvers like Glucose [3, 4] inspired by the earlier MiniSAT [25], via convenient interfaces like PySAT [30, 31], other fast SAT solvers such as Kissat [12], counterexample finders like Mace4 [43], and more general ATPs like Vampire [38] which can be used as solvers, all succeeded in finding useful magmas following this approach.
@@ 6.3. Canonization
As a conclusion for this section, we note that a very general strategy for building canonizers comes from the theory of rewrite systems, see e.g. Baader & Nipkow [7]. In that setting one defines rewriting as a transformation on words (or terms), and if this transformation is terminating and confluent (intuitively, rewrites cannot go on forever, or diverge forever), then one may simply pick the transformation which sends a term to its normal form as a canonizer.
@@ 6.4. Unique factorization
If (i) holds, then we have $\mathrm{x}\diamond\mathrm{y}=\mathrm{x}$ must hold in ${\mathcal{M}}_{X}/\sim_{\mathrm{E}}$, hence $\mathrm{E}854$ would imply $\mathrm{E}4$. However, the magma $\mathbb{Z}/2\mathbb{Z}$ with $x\diamond y=x(1-y)$ refutes this implication.
Similarly, if (iii) held, then $\mathrm{E}854$ would have to imply $\mathrm{E}10$, but this can be refuted by a finite magma with $11$ elements.